# Large Language Models Are Active Critics in NLG Evaluation

### v1

@@ 4.2 Self-Optimizing Scoring
In this module, the LLM assesses the model output $y_{i}$ based on the criteria set $C=\{c_{1},c_{2},\dots,c_{m}\}$ obtained from the task inference stage (§4.1). Specifically, for each input-output pair $(x_{i},y_{i})$, the LLM is prompted to estimate a score $\hat{r}_{ij}$ and a corresponding explanation $e_{ij}$ according to each criterion $c_{j}\in C$:
$\text{LLM}(x_{i},y_{i},f_{\text{McS-E}}(T,C,{\mathcal{D}}_{\text{demo}}))\to R_{i}=\{(\hat{r}_{ij},e_{ij}),~{}\forall c_{j}\in C\}$ (1)
where the output uses a JSON format, indicating a set of score-explanation pairs $R_{i}$ for all criteria in $C$ and ${\mathcal{D}}_{\text{demo}}$ is a set of demonstration examples randomly selected from the training set ${\mathcal{D}}_{\text{train}}$. This mechanism ensures that the evaluation is both quantitative and interpretive, offering insights into the rationale behind each score. The prompt template $f_{\text{McS-E}}(T,C,{\mathcal{D}}_{\text{demo}})$ is designed to enable scoring across multiple criteria simultaneously, accounting for the interconnections between them. This design allows for a fine-grained evaluation, where each criterion is treated both individually and in connection with the others, providing detailed explanations that enhance the transparency and interpretability of the scoring process.
@@
$\text{LLM}(x_{i},y_{i},f_{\text{OS-E}}(T,R_{i},{\mathcal{D}}_{\text{demo}}))\to\hat{r}_{i},e_{i}$ (2)
@@
Given the sensitivity of LLM performance to the few-shot demonstration examples ${\mathcal{D}}_{\text{demo}}$ in the prompt, we further propose an automatic prompt optimization strategy built upon DSPy (Khattab et al., 2023) to iteratively select the optimal ${\mathcal{D}}_{\text{demo}}$ to refine the prompts. Specifically, given two lists of overall quality scores across all examples in ${\mathcal{D}}_{\text{train}}$—one predicted by the LLM ($\hat{R}=\{\hat{r}_{i}\}_{i=1}^{N}$ from Eq. 2) and the other annotated by humans ($R$)—we design an objective function to maximize the correlation between these two score lists. To mitigate potential biases caused by relying on a single correlation measurement, we calculate the sum of three widely-used correlation coefficients: Pearson ($\gamma$), Spearman ($\rho$), and Kendall ($\tau$) with equal weights:
@@
$\Phi^{*}$ $=\arg\max_{\Phi}Q(\hat{R},R)$ (4)
where $\Phi=(T,C,{\mathcal{D}}_{\text{demo}})$ is an evaluation protocol consisting of a text description of the NLG task $T$, evaluation criteria $C$, and a few demonstration examples ${\mathcal{D}}_{\text{demo}}$ randomly selected from ${\mathcal{D}}_{\text{train}}$. To approximately solve the above maximization problem, we repeat $K$ time for the evaluations of Eq. 2 using different randomly sampled ${\mathcal{D}}_{\text{demo}}$, and select the best ${\mathcal{D}}_{\text{demo}}$ that maximizes $Q(\hat{R},R)$.
@@ 5 Experiment Settings
Following prior work (Zhong et al., 2022; Fu et al., 2024; Liu et al., 2023a), we employ four popularly-used benchmarks for meta-evaluation. These datasets cover diverse topics (e.g., politics, sports, restaurants, etc.) across four NLG tasks (i.e., summarization, dialogue generation, data-to-text generation, and storytelling), aiming to construct a robust testbed to access Active-Critic. The details of each benchmark are described below.
  • SummEval (Fabbri et al., 2021) consists of 1,600 machine-generated summaries based on CNN/DailyMail articles curated by  (Hermann et al., 2015). Each summary is annotated by both expert and layman judges on four aspects: coherence, consistency, fluency, and relevance. The dataset provides both aspect-specific scores and an overall quality score.
  • Topical-Chat (Mehri & Eskenazi, 2020) is a knowledge-grounded, open-domain conversation benchmark. It contains 60 conversations, each paired with 6 responses generated by different systems (2 from humans and 4 from machines). Each dialogue response is evaluated with human scores on overall quality, considering five aspects: naturalness, coherence, engagingness, groundedness, and understandability.
  • SFRES (Wen et al., 2015) is a data-to-text generation benchmark containing 1,181 instances. The task focuses on generating natural language utterances from structured restaurant information in San Francisco. The overall quality score is annotated by humans based on two aspects: informativeness and naturalness.
  • OpenMEVA (ROC) (Guan et al., 2021) collects open-ended commonsense stories generated by various models trained upon ROCStories corpus. It includes 1,000 data instances in total. Each story is rated by humans on its overall quality regarding three aspects: fluency, creativity, and coherence.
We standardize all benchmarks into a uniform format that includes: (1) the machine-generated texts to be evaluated, (2) the source input used by generation systems to produce these texts, and (3) the human scores assessing the generated outputs’ quality.
@@
We compare Active-Critic with a variety of state-of-the-art publicly accessible NLG evaluation methods. The baselines are grouped into two categories: (1) fine-tuned models including Auto-J  (Li et al., 2023) and UniEval (Zhong et al., 2022); and (2) prompting-based, train-free methods, including GPTScore (Fu et al., 2024), ExplainEval (Mahmoudi, 2023), and G-eval (Liu et al., 2023a). To ensure a fair comparison between the train-free baselines and Active-Critic, we use the same backbone LLM in comparisons. For GPTScore, we denote GPTScore-src to indicate the use of the src-hypo scoring type.
We focus on three correlation coefficients to assess the evaluation consistency between machine-based evaluators and humans: Pearson ($\gamma$(Mukaka, 2012), Spearman ($\rho$(Zar, 2005) and Kendall-Tau ($\tau$(Kendall, 1938).
Meta-evaluation
We establish Active-Critic using two widely adopted backbone models: one open-source LLM—Orca2-13b, and the other close-source LLM—GPT-3.5 (gpt-3.5-turbo-1106). In our preliminary study, we also tested Active-Critic on LLaMA2-13B. Since Orca2-13B outperformed LLaMA2-13B in achieving higher alignment with human judgments, we selected Orca2-13B for further analysis. Given the long-term accessibility and lower costs of open-source LLMs compared to closed-source models, we focus on Orca2-13B-based Active-Critic across four NLG tasks. For comparison, we built the GPT-3.5-based Active-Critic with a specific focus on the two most commonly used NLG tasks (SummEval and TopicalChat) in the meta-evaluation.
We focus on three variants of Active-Critic (AC) in the meta-evaluation. First, AC-Vanilla is a protocol that directly prompts the base LLM to score the overall quality of the input test case using self-optimizing scoring. Without providing or guiding the LLM to generate task-specific information, this variant explores the base LLM’s ability to make judgments based solely on its intrinsic understanding of few-shot graded exemplars. The second variant AC-Coarse, performs a coarse-grained, explainable evaluation by prompting the LLM to infer task-specific information (i.e., task definition and necessary criteria) and produce an overall score along with an explanation for each test case. This process considers all inferred criteria simultaneously during the self-optimizing scoring. Finally, AC-Fine provides a fine-grained, explainable evaluation. Similar to AC-Coarse, it begins with task inference, but during self-optimizing scoring, it assesses the input test case against each criterion individually, offering detailed explanations for each score. The overall quality score is then generated by combining the evaluations across all criteria. Details of the parameter settings and implementation are provided in Appendix B.
6 Results and Analysis
6.1 Overview of Active-Critic Performance
@@
Auto-J 0.1345 0.1457 0.1149 0.4681 0.459 0.3714 0.126 0.1022 0.0809 0.3896 0.3704 0.3065 0.2558
UniEval 0.5457 0.4914 0.3707 0.5133 0.5448 0.4134 0.2894 0.2499 0.1877 0.4501 0.4408 0.3119 0.4008
GPTScore-src 0.4043 0.3584 0.2696 0.2313 0.2437 0.1792 0.2062 0.121 0.1132 0.2283 0.2265 0.1534 0.2280
ExplainEval 0.5447 0.4916 0.3999 0.5542 0.5512 0.4476 0.1993 0.1181 0.0946 0.4809 0.4695 0.358 0.3924
Ours
AC-Vanilla 0.5494 0.486 0.3964 0.4932 0.4927 0.4055 0.2322 0.1712 0.1328 0.4023 0.4245 0.3209 0.3756
AC-Coarse 0.5386 0.5227 0.4156 0.611 0.6173 0.4845 0.3094 0.2663 0.199 0.4908 0.4962 0.3622 0.4428
AC-Fine 0.6301 0.5486 0.4299 0.6023 0.6214 0.4713 0.2915 0.2501 0.1906 0.5259 0.5363 0.4109 0.4591
@@
Active-Critic outperforms baseline evaluators across four distinct NLG tasks.
Table 1 displays the correlation results between unified evaluators and human judgments on four distinct NLG tasks. Overall, we observe that the coarse and fine variants of Active-Critic consistently exhibit a higher correlation with human judgments than baselines across three correlation coefficients in four NLG tasks. The fine-level evaluation generally outperforms the coarse variant. Among three variants, the vanilla Active-Critic performs the worst, even with lower average correlation compared to UniEval and ExplainEval. Our observations suggest that while the backbone LLM’s (i.e., Orca2-13B) intrinsic knowledge is limited for directly scoring data without task-specific inference, even with prompt optimization using a few examples, our proposed active-critic mechanism can effectively unlock the model’s potential to deeply digging into the example data for evaluation. Moreover, this mechanism enhances the model’s ability to assess data quality with greater alignment to human judgments. Notably, guiding the model to evaluate each inferred criterion individually leads to better final decisions than asking it to directly provide an overall score.
SummEval TopicalChat
$\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$
G-eval 0.4687 0.4504 0.3745 0.5427 0.5597 0.4501
Ours
AC-Coarse 0.6569 0.5368 0.4178 0.6425 0.6171 0.4855
AC-Fine 0.653 0.6016 0.4745 0.6718 0.6703 0.5156
Table 2: Correlation between GPT-3.5-based evaluators and human judgments on instance-level overall quality for SummEval and TopicalChat. We compare Pearson ($\gamma$), Spearman ($\rho$) and Kendall-Tau ($\tau$) correlation, respectively. The best performance per indicator is highlighted in bold, and the second-highest results are underlined. We implemented and tested all the methods with p-value $<$ 0.05.
Active-Critic can achieve further improvements with a stronger backbone LLM.
Table 2 displays the results of Active-Critic built on top of GPT-3.5, with an emphasis on the task of SummEval and TopicalChat.
Comparatively, Active-Critic built on GPT-3.5 shows significantly higher correlations with human judgments than the variant built on Orca-13B, indicating that a stronger backbone model enhances the effectiveness of our evaluation protocol. Further comparison with the state-of-the-art GPT-3.5-based baseline reveals that Active-Critic consistently outperforms it. As with the Orca-13B-based Active-Critic (see Table 2), the fine-level variant surpasses the coarse one, further suggesting that multi-criteria scoring and explanations help the backbone LLM conduct a more in-depth analysis of instance quality and make better final decisions.
6.2 Explainability Analysis
Are the generated explanations helpful to Active-Critic’s scoring?
To assess the impact of explanations generated by Active-Critic, we compared our protocol’s performance with versus without explanations, at both coarse and fine levels of evaluations. Figure 2 shows the results based on the Kendall-Tau correlation. We also provide the results of Pearson and Spearman correlation in Appendix D.
Human Evaluation.
Dimension Clarity Relevance Score Consistency Accuracy Aspect-to-Overall Alignment Differentiability Overall

### v2

@@ 4.2 Scoring Alignment
In this module, we use the LLM to assess the model output $y_{i}$ based on the criteria set $\hat{C}=\{\hat{c}_{1},\hat{c}_{2},\dots,\hat{c}_{m}\}$ obtained from the task inference stage (§4.1). Specifically, for each input-output pair $(x_{i},y_{i})$, the LLM is instructed to estimate a score $\hat{r}_{ij}$ and a corresponding explanation $e_{ij}$ according to each criterion $\hat{c}_{j}\in\hat{C}$:
$\text{LLM}(x_{i},y_{i},f_{\text{McS-E}}(T,C,{\mathcal{D}}_{\text{demo}}))\to\hat{R}_{i}$ (1)
$\hat{R}_{i}=\{(\hat{r}_{ij},\hat{e}_{ij}),~{}\forall\hat{c}_{j}\in\hat{C}\}$ (2)
where the output uses a JSON format, indicating a set of score-explanation pairs $\hat{R}_{i}$ for all criteria in $\hat{C}$ and ${\mathcal{D}}_{\text{demo}}$ is a set of demonstration examples randomly selected from the training set ${\mathcal{D}}_{\text{train}}$. This mechanism ensures that the evaluation is both quantitative and interpretable, offering insights into the rationale behind each score. The prompt template $f_{\text{McS-E}}(T,C,{\mathcal{D}}_{\text{demo}})$ is designed to enable scoring across multiple criteria simultaneously, accounting for the interconnections between them. This design enables a fine-grained evaluation, where each criterion is treated both individually and in connection with the others, providing detailed explanations that enhance the interpretability of the scoring process.
@@
$\text{LLM}(x_{i},y_{i},f_{\text{OS-E}}(T,\hat{R}_{i},{\mathcal{D}}_{\text{demo}}))\to\hat{r}_{i},\hat{e}_{i}$ (3)
@@
Given the sensitivity of LLMs’ in-context prediction performance to the few-shot examples ${\mathcal{D}}_{\text{demo}}$ in the prompt, we further propose an automatic prompt optimization strategy built upon DSPy (Khattab et al., 2023) to iteratively select the optimal ${\mathcal{D}}_{\text{demo}}^{*}$ to refine the prompts. Specifically, given two lists of overall quality scores across all examples in ${\mathcal{D}}_{\text{train}}$—one predicted by the LLM, i.e., $\hat{r}=[\hat{r}_{1},\dots,\hat{r}_{N}]$ from Eq. (3), and the other annotated by humans, i.e., $r=[r_{1},\dots,r_{N}]$—we design an objective function to maximize the correlation between these two score lists. To mitigate potential biases caused by relying on a single correlation measurement, we calculate the sum of three widely-used correlation coefficients: Pearson ($\gamma$), Spearman ($\rho$), and Kendall ($\tau$) with equal weights:
@@
${\mathcal{D}}_{\text{demo}}^{*}$ $=\arg\max_{{\mathcal{D}}_{\text{demo}}\subset{\mathcal{D}}}{\mathcal{Q}}(\hat{r},r)$ (5)
where ${\mathcal{D}}_{\text{demo}}$ is the optimal few-shot demonstration examples ${\mathcal{D}}_{\text{demo}}$ selected from ${\mathcal{D}}_{\text{train}}$. To approximately solve the above maximization problem, we repeat $K$ time for the evaluations of Eq. (3) using different randomly sampled ${\mathcal{D}}_{\text{demo}}$, and select the best ${\mathcal{D}}_{\text{demo}}^{*}$ that maximizes ${\mathcal{Q}}(\hat{r},r)$.
@@ 5 Experiment Settings
Following prior work (Zhong et al., 2022; Fu et al., 2024; Liu et al., 2023a), we evaluate our method on four popularly-used benchmarks. These datasets cover diverse topics (e.g., politics, sports, restaurants, etc.) across four NLG tasks (i.e., summarization, dialogue generation, data-to-text generation, and storytelling), aiming to construct a robust testbed to access Active-Critic. The details of each benchmark are described below.
  • SummEval (Fabbri et al., 2021): 1,600 machine-generated summaries of CNN/DailyMail articles were rated by both expert and layman judges on coherence, consistency, fluency, relevance, and overall quality.
  • Topical-Chat (Mehri and Eskenazi, 2020): A knowledge-grounded, open-domain dialogue dataset consisting of 60 conversations, each paired with 6 responses (2 by humans and 4 by machines). Responses are human-evaluated on overall quality across five dimensions: naturalness, coherence, engagingness, groundedness, and understandability.
  • SFRES (Wen et al., 2015): A data-to-text generation benchmark with 1,181 instances, focusing on generating free-text utterances from structured restaurant information. Annotators rated the overall quality of each instance based on informativeness and naturalness.
  • OpenMEVA (ROC) (Guan et al., 2021): 1,000 open-ended commonsense stories generated by various models trained upon the ROCStories corpus. Annotators rate each story based on fluency, creativity, and coherence.
We standardize all benchmarks into a uniform format that includes: (1) the machine-generated responses for evaluation, (2) the source input used by the generation systems for response generation, and (3) the human scores assessing response quality.
@@
We compare Active-Critic with a variety of state-of-the-art publicly accessible NLG evaluation methods. The baselines are grouped into two categories: (1) fine-tuning-based methods including Auto-J  (Li et al., 2023), UniEval (Zhong et al., 2022), InstructScore (Xu et al., 2023) and TIGERScore (Jiang et al., 2023); and (2) prompting-based methods, including GPTScore (Fu et al., 2024), G-eval (Liu et al., 2023a) and four selected base LLMs under the zero-shot manner, implemented following (Mahmoudi, 2023). Following prior work (Fu et al., 2024; Jiang et al., 2023), we use GPTScore-src to refer to the source-hypothesis scoring type, which is a reference-free evaluation method.
Regarding metrics, we use three correlation coefficients to assess the evaluation consistency between machine-based evaluators and humans: Pearson ($\gamma$(Mukaka, 2012), Spearman ($\rho$(Zar, 2005) and Kendall-Tau ($\tau$(Kendall, 1938).
@@
Fine-tuning LLM
InstructScore 0.3496 0.2703 0.203 0.2691 0.2774 0.2423 0.2039 0.1502 0.133 0.2936 0.2772 0.1658 0.2363
Auto-J 0.1345 0.1457 0.1149 0.4681 0.459 0.3714 0.1315 0.1053 0.0869 0.3896 0.3704 0.3065 0.257
TIGERScore 0.458 0.3694 0.2937 0.3785 0.4401 0.3458 0.1898 0.1246 0.1075 0.451 0.4413 0.3356 0.3279
UniEval 0.5457 0.4914 0.3707 0.5133 0.5448 0.4134 0.3247 0.2791 0.2081 0.4501 0.4408 0.3119 0.4078
Prompting Open-source LLM
GPTScore-src (FLAN-T5) 0.4043 0.3584 0.2696 0.2313 0.2437 0.1792 0.2819 0.2082 0.1618 0.2283 0.2265 0.1534 0.2456
Zero-shot (LLaMA3-8B) 0.4104 0.3857 0.2809 0.5197 0.5242 0.4018 0.2138 0.196 0.152 0.4141 0.3676 0.2808 0.3456
Zero-shot (Orca2-13B) 0.5447 0.4916 0.3999 0.5542 0.5512 0.4476 0.3068 0.23 0.1842 0.4809 0.4695 0.358 0.4182
Ours:
AC-Coarse (LLaMA3-8B) 0.5307 0.4972 0.3958 0.4873 0.5246 0.4259 0.1853 0.1594 0.1451 0.4394 0.4403 0.3477 0.3816
AC-Fine (LLaMA3-8B) 0.5334 0.502 0.401 0.5321 0.5379 0.4045 0.2265 0.2245 0.169 0.4506 0.4436 0.3625 0.399
[…]
AC-Fine (Orca2-13B) 0.6301 0.5486 0.4299 0.6023 0.6214 0.4713 0.324 0.2834 0.2289 0.5259 0.5363 0.4109 0.4677
Prompting Close-source LLM
G-eval (GPT-3.5) 0.4687 0.4504 0.3745 0.5427 0.5597 0.4501 0.2464 0.1956 0.1591 0.362 0.3408 0.1982 0.3624
Zero-shot (GPT-3.5) 0.453 0.385 0.292 0.5503 0.5436 0.4231 0.2823 0.2274 0.1828 0.4229 0.397 0.3 0.3716
Zero-shot (GPT-4) 0.5943 0.5038 0.4055 0.6659 0.656 0.4937 0.3301 0.2823 0.2284 0.5627 0.4928 0.3777 0.4661
Ours:
AC-Coarse (GPT-3.5) 0.6569 0.5368 0.4178 0.6425 0.6171 0.4855 0.3585 0.2846 0.2374 0.4185 0.3766 0.2981 0.4442
AC-Fine (GPT-3.5) 0.653 0.6016 0.4745 0.6718 0.6703 0.5156 0.3616 0.2833 0.2342 0.4693 0.4527 0.3442 0.4777
[…]
AC-Fine (GPT-4) 0.6926 0.5723 0.462 0.7789 0.7753 0.6212 0.363 0.2809 0.236 0.5877 0.5581 0.4249 0.5294
@@
Meta-evaluation
We establish Active-Critic using four widely adopted backbone models: two open-source LLMs (Orca2-13B and LLaMA3-8B) and two closed-source LLMs (GPT-3.5 and GPT-4) [^2] We used GPT-3.5-turbo-1106 and gpt-4-turbo version for the experiments. across four diverse NLG tasks. We test two variants of Active-Critic (AC) in this study: (1) AC-Coarse performs a coarse-grained, explainable evaluation by prompting the LLM to infer task-related information and directly produce an overall score along with an explanation for each test case. This process considers all inferred criteria at once during scoring alignment. (2) AC-Fine provides a fine-grained, explainable evaluation. Similar to AC-Coarse, it begins with task inference, but during scoring alignment, it assesses the input test case against each criterion individually, offering detailed explanations for each score. The overall quality score is then generated by combining the evaluations across all criteria. Appendix B provides the details of implementation.
6 Results and Analysis
SummEval TopicalChat SFRES OpenMEVA (ROC) Average
$\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$ $\gamma$ $\rho$ $\tau$
Ours (AC-Fine) 0.6301 0.5486 0.4299 0.6023 0.6214 0.4713 0.324 0.2834 0.2289 0.5259 0.5363 0.4109 0.4677
w/o Task Description 0.5825 0.4826 0.3552 0.4949 0.5057 0.4211 0.2683 0.2017 0.168 0.3846 0.3802 0.2918 0.3781
w/o Criteria Definition 0.5726 0.522 0.4062 0.5533 0.5368 0.4451 0.293 0.2715 0.1907 0.4176 0.4237 0.326 0.4132
w/o McS-E 0.5386 0.5227 0.4156 0.611 0.6173 0.4845 0.3612 0.2981 0.2393 0.4908 0.4962 0.3622 0.4531
w/o OS-E 0.6106 0.5129 0.3908 0.5639 0.5615 0.4464 0.3165 0.2405 0.1899 0.509 0.4931 0.3632 0.4332
Table 2: Ablation study of key modules in Active-Critic.
6.1 How well does Active-Critic perform?
Table 5 displays the correlation between unified evaluators and human judgments. Overall, Active-Critic noticeably outperforms the corresponding prompting-based baselines and state-of-the-art fine-tuning-based evaluators, where our variants built on Orca2-13B and GPT-4 achieve the highest correlation in the methods using open- and close-source LLM, respectively. Comparing two variants of Active-Critic per LLM, we find that the fine-level variant consistently achieves higher alignment with human scores, outperforming the coarse one by $\sim$2in average correlation. These results show that our approach can effectively enhance LLMs’ potential to capture human-centric assessment nuances in diverse scenarios and make more human-aligned judgments. Moreover, prompting the LLM to assess each criterion individually and then aggregate the scores benefits Active-Critic’s decision-making.
We also validate the generalizability of Active-Critic’s self-inferred evaluation prompts on unseen data within a similar NLG scenario. We use the prompts generated by Active-Critic on SummEVAL examples to evaluate responses from the Newsroom (Grusky et al., 2018) benchmark, similarly focusing on news article summarization but using diverse strategies. As shown in Appendix C, our approach outperforms state-of-the-art baselines, demonstrating strong generalization.
Further examining our approach’s stability across base LLMs, we observe that Active-Critic consistently achieves a noticeable improvement, with an average gain of $\sim$6.8correlation over the zero-shot baseline for each base model. This indicates its effectiveness, regardless of the chosen base LLM. Although Active-Critic generally obtains greater enhancements when employing a stronger base LLM, it is noteworthy that Active-Critic built on Orca2-13B performs comparably to its GPT-4 counterpart on SFRES and OpenMEVA (ROC). Considering the computational cost and evaluator performance, we primarily focus on the ORCA2-based AC-Fine for further analysis.
6.2 Ablation Study
Dependence on Human-scored Data.
To examine the impact of labeled data size on Active-Critic’s performance, we varied the size of the feed examples from 5-shot to 5%, 15%, and the full 25% of each benchmark. Figure 2 shows the results. While Active-Critic improves as the labeled data size increases, it can achieve a decent correlation with human evaluators using as few as five human-rated examples. Among four tasks, Active-Critic is more sensitive to labeled data size in TopicalChat and SummEval than in OpenMeva and SFRES. The former two benchmarks involve longer contexts and diverse topics, while the latter focus on specific topics with shorter contexts, making the first two tasks more complex. Our observations suggest that Active-Critic requires more labeled data for evaluating complex NLG tasks compared to simpler ones.
Figure 2: Average correlation between Orca2-based Active-Critic and human judgments with varying label sizes. Results for each correlation coefficient are provided in Appendix D
Impact of Optimization.
We compare Active-Critic’s performance by removing its dynamic prompt optimization for scoring and, furthermore, eliminating mini-batch iterations during task inference. As shown in Figure 3, there is a drop in Active-Critic’s performance when removing scoring prompt optimization, with a further decline when only using a single mini-batch of labeled data for task inference, suggesting that both strategies contribute to Active-Critic for making optimal decisions. Interestingly, the Active-Critic shows greater sensitivity to scoring optimization in the fine-level evaluation of SummEval and the coarse-level evaluation of SFRES, indicating that this component plays a more significant role in these specific evaluation scenarios. In contrast, the influence of mini-batch iterations for task inference is minimal in SummEval, suggesting that Active-Critic can effectively infer the target evaluation task in this setting with limited training data.
Figure 3: Impact of prompt optimization on scoring and mini-batch iterations on task inference (Kendall-Tau %). See Appendix E for Pearson and Spearman results.
Coherence: The degree to which the summary flows logically and cohesively, with clear connections between the main points. Conciseness: The ability of the summary to convey all necessary information in a succinct and efficient manner. Coverage: The extent to which the summary captures the main events and details from the source text without omitting crucial information. Accuracy: The faithfulness of the summary to accurately reflect the main points and details of the source text. Fluency: The readability and naturalness of the language used in the summary, with smooth transitions between ideas and paragraphs. Relevance: The relevance of the summary to the main topic and the inclusion of only pertinent information from the source text. Clarity: The clarity and comprehensibility of the summary, with clear and precise language used to convey the main points. Engagement: The ability of the summary to captivate and engage the reader, drawing them into the main events and details effectively. Coherence: the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic. Consistency: the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Fluency: the summary should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read. Relevance: The summary should include only important information from the source document.
(a) AC-Fine (b) Human
Table 3: An illustrative example of the generated evaluation criteria on SummEval, either generated by an Active-Critic (a) or predefined by humans (b). The highlighted text in blue are additional criteria generated by the machine compared to the human-defined ones.
Module Contribution.
Table 2 shows the individual contribution of each module in Active-Critic. Note that the variant w/o criteria inference uses the original predefined criteria from each benchmark for further computation. In the variant w/o OS-E, we calculated the overall quality score per test case by averaging the multiple criteria-specific scores generated by McS-E. We find that removing task inference modules leads to a more substantial performance drop compared to removing scoring modules, especially when the LLM is not asked to infer the task description (resulting in a $\sim$9decrease on average). Our findings suggest that guiding LLMs to understand what to evaluate contributes more to Active-Critic’s effectiveness than other modules. Additionally, the larger performance drop in the variant w/o OS-E, compared to the one w/o McS-E, indicates that the LLM-generated overall quality score contributes more meaningfully than simply averaging the criteria-specific scores.
6.3 Qualitative Analysis of Active-Critic
LLM-inferred Criteria Analysis.
Moving forward from quantitative analysis, we examine the LLM-inferred criteria in depth. Table 3 shows an illustrative comparison between the criteria generated by Active-Critic and those pre-defined by humans in SummEval. We find that our approach incorporates more nuanced criteria (i.e., “clarity”, “conciseness”, “coverage”, and “engagement”) beyond the four pre-defined aspects. Moreover, each criterion is paired with a clear definition to specify its distinct characteristics. For example, the human-defined “coherence” starts with a high-level description like “well-structured and well-organized”, while the LLM’s definition tends to be more concrete, e.g., “the summary flows logically”.
Dimension Clarity Relevance Score Consistency Accuracy